OpenStudy (babyslapmafro):

How do you find the anti-derivative of e^(-2x)?

OpenStudy (anonymous):

\[\frac{ d }{ dx }(e^{-2x})=-2x*e^{-2x}\]

OpenStudy (babyslapmafro):

I'm trying to find the anti-derivative

OpenStudy (babyslapmafro):

And the derivative you found is incorrect

OpenStudy (anonymous):

then integrate it

OpenStudy (anonymous):

sorry i read it derivative

jimthompson5910 (jim_thompson5910):

Let u = -2x so du/dx = -2 du = -2*dx dx = -du/2

OpenStudy (babyslapmafro):

ok, thanks...